# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 23 $\overline{2}\overline{3}4,4\overline{2},\ \overline{3}\overline{4},\overline{4}$ 3,1, 5,4 0,1, 3,2 9
Next, we consider a different formula for the restriction $\Phi|_{\mathsf{X}_{1}}$ of $\Phi$ to $\mathsf{X}_{1}$. For $x\in\mathsf{X}_{1}$, let $y({\boldsymbol{u}})=\varphi(x)\in Y_{1}$, so all crossings in $y({\boldsymbol{u}})$ are $2$-fold. By Lemma 5.8, the Wang patches $\{w_{\boldsymbol{n}}(x)\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ in $x$ are the duals of the grid patches $\{g_{\boldsymbol{n}}({\boldsymbol{u}})\}_{{\boldsymbol{n}}\in\mathbb{Z}^{2}}$ in $y({\boldsymbol{u}})$, namely, $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Define $\theta:Y_{1}\to W_{Aq_{\rm p}}$ to be the map that replace each grid patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})$ with the Wang tile $\Theta(g_{\boldsymbol{n}}({\boldsymbol{u}})):=AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})$, noting as above, that both $g_{\boldsymbol{n}}({\boldsymbol{u}})$ and $AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})$ are supported on $R_{\boldsymbol{n}}$. In particular, $\theta$ is a local map (see Remark 3.4), so it is uniformly continuous, and satisfies $\theta(T^{\boldsymbol{t}}y({\boldsymbol{u}}))=T^{\boldsymbol{t}}(\theta(y({\boldsymbol{u}})))$. Clearly $\Phi|_{\mathsf{X}_{1}}(x)=\theta(\varphi(x))$. Then $\Phi|_{\mathsf{X}_{1}}$ is uniformly on the dense set $\mathsf{X}_{1}\subseteq\mathsf{X}$, and satisfies $\Phi|_{\mathsf{X}_{1}}(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi|_{\mathsf{X}_{1}}(x))$. So $\Phi$ is its unique continuous extension, satisfying $\Phi(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi(x))$.
It remains to show $\Phi$ is a bijection. Lemma 6.7 shows there is a 1:1 correspondence between $T$ orbits on $\mathsf{X}$ and $T$ orbits on $W_{Aq_{\rm p}}$. But since $T$ acts freely on both $\mathsf{X}$ and $W_{Aq_{\rm p}}$, $\Phi(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi(x))$ shows that $\Phi$ is bijective on each orbit. This implies $\Phi$ is a bijection. ∎

### v2

@@ 23 $\overline{2}\overline{3}4,4\overline{2},\ \overline{3}\overline{4},\overline{4}$ 3,1, 5,4 0,1, 3,2 9
Next, we consider a different formula for the restriction $\Phi|_{\mathsf{X}_{1}}$ of $\Phi$ to $\mathsf{X}_{1}$. For $x\in\mathsf{X}_{1}$, let $y({\boldsymbol{u}})=\varphi(x)\in Y_{1}$. By Lemma 5.8, $w_{\boldsymbol{n}}(x)=g_{\boldsymbol{n}}({\boldsymbol{u}})^{*}$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{2}$. Define $\theta:Y_{1}\to W_{Aq_{\rm p}}$ to be the map that replaces each grid patch $g_{\boldsymbol{n}}({\boldsymbol{u}})$ in $y({\boldsymbol{u}})$ with the Wang tile $\Theta(g_{\boldsymbol{n}}({\boldsymbol{u}})):=AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})$, noting that both $g_{\boldsymbol{n}}({\boldsymbol{u}})$ and $AQ^{a_{\boldsymbol{n}}}+A({\boldsymbol{n}}-{\boldsymbol{u}}^{\prime})$ are supported on $R_{\boldsymbol{n}}$. In particular, $\theta$ is a local map in the sense of Definition 3.5, so it is uniformly continuous, and satisfies $\theta(T^{\boldsymbol{t}}y({\boldsymbol{u}}))=T^{\boldsymbol{t}}(\theta(y({\boldsymbol{u}})))$. Clearly $\theta(\varphi(x))=\Phi|_{\mathsf{X}_{1}}(x)$ (from the first part), so $\Phi$ is uniformly continuous on $\mathsf{X}_{1}$, and satisfies $\Phi|_{\mathsf{X}_{1}}(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi|_{\mathsf{X}_{1}}(x))$. It follows that $\Phi$ is the unique continuous extension of $\theta(\varphi(x))$, and thus satisfies $\Phi(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi(x))$.
It remains to show $\Phi$ is a bijection. Lemma 6.7 shows there is a 1:1 correspondence between $T$ orbits on $\mathsf{X}$ and $T$ orbits on $W_{Aq_{\rm p}}$. But since $T$ acts freely on both $\mathsf{X}$ and $W_{Aq_{\rm p}}$, the fact that $\Phi(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}(\Phi(x))$ shows that $\Phi$ is bijective on each orbit as well. This implies $\Phi$ is a bijection. ∎